# WikiGap: Promoting Epistemic Equity by Surfacing Knowledge Gaps Between English Wikipedia and other Language Editions

### v3

@@ 2.1. English Wikipedia is Considered the Default, yet Information Gaps with Other Editions Persist
2.2. Prior Work Targeting Cross-Lingual Knowledge Dissemination
Researchers have taken notice of the complementary nature of Wikipedia’s language editions, proposing various technical solutions to exposing users to information from different editions.
Towards a Monolithic Language Edition.
Differences between language editions that stem from socio-geographical interests have often been perceived as rather cross-linguistic disparities. Underpinning the framing of disparities is that there is a problem to be solved. Consider, for example, this quote from Duh et al. (2013), who propose to automatically construct a unified version of an article across languages:
Information disparity is a major challenge with multi-lingual document collections. When documents are dynamically updated in a distributed fashion, information content among different language editions may gradually diverge.
This problem formulation is not uncommon, even serving as the locus of a prominent shared task (Negri et al., 2013). Adar et al. (2009) take a similar approach and propose to unify information boxes across language editions, pursuing consolidation in Wikipedia infoboxes. Conversely, aligned with prior work in understanding cross-linguistic heterogeneity across Wikipedia’s language editions (e.g., Bao et al., 2012; Sen et al., 2015), we take the view that cultural positionality is one to be highlighted rather than a problem that needs to be solved via consolidation into a monolithic knowledge base, where different languages serve as mere cryptographic encodings over raw facts.
Wikipedia’s Interlanguage Links (ILL)
Currently, Wikipedia supports accessing different language editions of a topic through Interlanguage Links, available via a dropdown box on each article. While these links ostensibly provide readers with access to complementary information across editions, the interface gives little indication that articles may differ substantially in content. Instead, the hyperlinks imply only a change in linguistic encoding, similar to using Google Translate, and do not signal the significant differences in composition and curation across editions. Consistent with this design, our user studies (Section 5) show that many readers assume language editions share a single underlying knowledge base.
As Samir et al. (2024) identified through their InfoGap method (Sec. 3.2), English Wikipedia lacks facts from other language editions that may have a broader appeal. For example, Apple CEO Tim Cook’s association with the Russia-Ukraine war is only mentioned on Russian Wikipedia, but is conceivably noteworthy to the wider array of readers that engage with English Wikipedia.
@@ 2.2. Prior Work Targeting Cross-Lingual Knowledge Dissemination
Prior Multilingual Interface Designs.
There have been attempts to go beyond both a unified language edition as well as Wikipedia’s ILL and develop interfaces to surface information from various language editions. One such system is Manypedia (Massa and Scrinzi, 2012), an interface providing a side-by-side view of an article from two language editions, translated into a common language using Google Translate. This interface, however, does not directly surface information gaps and inconsistencies between the two language editions. In contrast, the Chrome extension WikiCompare (Roy et al., 2022) identifies differences in topic coverage. In the typical scenario, a user reading an article in their native language (e.g., Hindi) will see highlighted phrases representing topics for which articles exist in another language edition (e.g., English) but not in the current language edition. Clicking on one of these links directs the user to the section of the English edition that discusses the topic. In our work, we go beyond identifying differences in topic coverage to pinpointing specific facts covered in one language but not another.
The most pathbreaking of these interfaces is Omnipedia (Bao et al., 2012), which enables readers to get an overarching sense of how different language editions’ editor communities conceptualize a concept. While they similarly rely on anchor links as WikiCompare does, their design enables visualizing several different language editions simultaneously and succinctly. This innovative redesign, however, would require users to completely change how they interact with Wikipedia. The Omnipedia workflow requires that readers start by entering a concept in their search bar. This starting point, however, contrasts with how most readers engage with Wikipedia, through information brokering search engines (Vincent and Hecht, 2021). As such, this innovative redesign is unlikely to meet widespread adoption.
In the next section, we describe WikiGap, a system powered by InfoGap (Samir et al., 2024), that supplements English Wikipedia pages with relevant facts from other language editions, providing the reader with more information and, in doing so, raising awareness of unique content in other language editions. As such, it dispels the English-as-superset conceptual model (Hecht, 2013) that, as our user studies demonstrate, remains remarkably prevalent. Unlike all of the aforementioned interfaces, our approach is not restricted to surfacing only differences in hyperlinks between language editions.
@@ 3. WikiGap System
The goal of WikiGap is to challenge the epistemic superiority of English Wikipedia, by increasing interaction with complementary content present in smaller language editions. Guided by this goal, we combine a computational pipeline that identifies sentence-level gaps across languages with a user interface designed to surface this content through in-situ annotations and structured fact cards. We begin with outlining the design requirements informed by user needs and foundational theories that shaped the WikiGap interface (§3.1). We then step behind the scenes to introduce InfoGap (Samir et al., 2024), which we use to detect language-exclusive facts lacking in English Wikipedia (§3.2). Furthermore, we describe how we extended and adapted InfoGap to detect and deliver language-exclusive facts, including how we filter, translate, and bind them to UI components at runtime (§3.3). We conclude this section by outlining the system implementation details (§3.4).
@@ 3.1.1. Design Requirements Informed by User Needs and Theory
When it come to start designing the interface, we combined our findings from the exploratory interview with established theories that explain how people seek, process, and comprehend information while reading to define a set of design requirements for WikiGap. These consist of the overarching goal, helping readers seamlessly access cross-lingual content while maintaining the familiar experience of browsing English Wikipedia. The findings from the exploratory interview, requirements, and design elements are summarized in Figure 2 and detailed below.
@@
We conducted an exploratory interview with four Computer Science graduate students (2 women, 2 men), all regular Wikipedia users; three spoke an additional language (Korean, Russian, or Hindi). Each 30-minute session followed an unstructured, conversational format guided by open-ended questions about their multilingual Wikipedia usage, perceptions of the Wikipedia ILL system, and ideas for improving cross-lingual content presentation. Participants reviewed an English Wikipedia article on mooncake–a topic unfamiliar to them–and reflected on how multilingual facts could be surfaced in this context. They proposed interface features to better support multilingual reading habits. We took notes during the sessions and synthesized them using thematic coding to derive the following design requirements for WikiGap.
@@
Participants expressed a strong desire to access multilingual information without breaking their reading experience. They reported rarely using Wikipedia’s ILLs (F1), noting the disorienting nature of full-page switches: “Even if I click the other language, I don’t know where to look for the thing I care about. It’s like starting over.” Such page jumps impose high navigation costs (Pirolli and Card, 1999), requiring users to reorient to a new layout and language, which discourages exploration.
@@
R2: Ensure clear language attribution and traceability of multilingual facts to promote creditability.
In our interviews, a recurring concern among participants was the trustworthiness of information surfaced from other language editions–particularly when such facts were not present in the English version. Participants emphasized that they wanted to know “where the information comes from,“ including which language edition it originated from and the ability to read it in its original context (F3), because it is crucial for accessing information’s credibility. This requirement highlights the importance of traceability and attribution in multilingual information access. Readers must be able to understand both the linguistic source and the original context of each surfaced fact. This entails more than just showing a translation, but rather also requires visual and interactive mechanisms that reinforce provenance and credibility.
@@
User feedback revealed a range of interaction preferences: users want flexible access modes–both passive discovery and targeted retrieval (F4). Different participants had different expectations for how cross-lingual facts should be surfaced. Some favored passive discovery, where the system automatically highlights interesting differences without requiring user input. Others preferred more active control, such as being able to search for specific content or filter information by language or topic. They discussed concern over potential information overload if too many facts were presented at once.
@@ 3.1.2. Core Design Elements
To address R1 (non-disruptive presentation) and R2 (traceability), we implemented a subtle underlining system that signals the presence of additional information from other language editions. Each marker is country-color coded and appears as a patterned underline within the English article. When clicked, the side panel retrieves the fact relevant to that sentence, where the extension identifies complementary or missing information from another language edition. The underline then becomes more prominent, providing interaction feedback and inviting further exploration. Each language is associated with a distinct color – red for Chinese, blue for French, and green for Russian. This color coding is applied across both in-text highlights and corresponding elements in the side panel, reinforcing source attribution as emphasized in R2. As we elaborate in Sec. 3.3.2, only significant knowledge gaps identified by InfoGap are highlighted, rather than minor variations or differences in wording. This selective highlighting prevents information overload and ensures that users’ attention is drawn to substantive cross-lingual differences.
@@
To fulfill R1 and R2, we developed a margin-anchored sidebar that displays the facts, translated to English and organized by their source language. When a user interacts with an in-text highlight, the corresponding fact card is revealed in the sidebar, which appears on the right side of the article. The sidebar is collapsible and can be pinned for extended browsing or hidden entirely to maintain a minimalist reading view. Each language section presents its facts in dedicated groups. Each fact is shown in English, accompanied by attribution information, including the source language tag and link to the original article. The close spatial relationship between in-text cues and sidebar content supports R1’s emphasis on minimizing reading disruption.
@@
At the heart of WikiGap’s cross-lingual surfacing system is the fact card interface, which presents individual facts from other language editions in a compact, standardized format. Each fact card includes: (1) the target fact, translated into English, (2) a color-coded badge indicating the source language, and (3) a hyperlink labeled “View on [Language] Wikipedia” that opens the source article in a new tab with the original sentence highlighted. It mainly address R2 (transparent attribution), because it offers enough context to understand the fact at a glance, while providing a direct pathway to its original source for users who wish to verify or explore further.
@@ 3.2. The InfoGap Pipeline
WikiGap relies on InfoGap, a previously developed computational pipeline introduced by Samir et al. (Samir et al., 2024), which detects factual misalignments between Wikipedia articles across languages. This subsection introduces the core components of the original InfoGap system, which we use to identify sentence-level knowledge gaps between English and other language editions. In the following section, we will describe how we extended this pipeline to support additional languages and adapted its outputs for integration into the WikiGap interface.
@@ 3.3. Extending and Integrating InfoGap for WikiGap
3.3.1. Enabling Chinese language support.
The original InfoGap pipeline was developed for detecting cross-lingual gaps in biographical articles between English and either Russian or French. To generalize the method to a new domain and an additional language, we incorporated support for Chinese into the fact decomposition and verification modules. This involved custom preprocessing for tokenization, crafting prompts, and conducting human evaluation on Chinese texts to confirm that InfoGap performs comparably to its performance on the original languages studied by Samir et al. (2024).
@@ 4. Evaluation Methods
To validate our design, we conducted a mixed-methods user evaluation study, combining quantitative performance measures and qualitative feedback, which are summarized in Table 1 and detailed below. We aim to understand WikiGap’s usability, its impact on knowledge gain, and user perceptions about the system and the cross‐lingual content. [^3] The study protocol received approval from our institutional REB.
@@ 4.1. Participants
The study followed a within-subject design, in which each participant was assigned two out of five topics (Sec. 4.3). Each participant read one article using the WikiGap extension and another article using the default Wikipedia interface (with ILLs) as a control. The assignment of treatment and control conditions was randomized for each participant.
@@ 4.3. Materials and Topic Assignment
Topic assignment was designed to minimize cultural confounds and familiarity‑driven performance differences. Specifically, the food topic chosen for each participant could not be strongly linked to the participant’s self‑reported cultural or ethnic background. To reduce individual-performance bias, each topic was assigned in every condition to exactly four different participants.
@@ 5. Findings
We report findings from our mixed-method evaluation of WikiGap. We present five key findings that drive our understanding towards the user perceptions of multilingual knowledge engagement. These findings highlight how WikiGap reshapes users’ awareness of content asymmetries, supports multilingual exploration, and triggers additional reflections on the role of provenance in a multilingual and LLM-driven knowledge landscape.
①Users Strongly Preferred WikiGap over the Default ILL (D1, D2, D3)
We evaluated user preference using the System Usability Scale (SUS). The average SUS scores for each condition are presented in Figure 6, along with general SUS cutoff scores for fair, good, and excellent usability (Bangor et al., 2009). Overall, WikiGap achieved a substantially higher usability score compared to the control. WikiGap falls within the “excellent” usability category, while the default Wikipedia page with ILL scored below the “fair” threshold. A paired-samples t-test revealed a significant difference in scores between the WikiGap condition (M = 82.5, SD = 10.9) and the control condition (M = 48.2, SD = 18.6), t(20) = 6.60, p ¡ .001.
Figure 6. Box plot showing the System Usability Scale (SUS) scores for each condition. Blue represent the control condition (no WikiGap), and orange represent the treatment condition (with WikiGap). Horizontal dashed lines represent standard usability benchmarks in varying shades of green: light green for Fair usability (SUS ¿ 51), medium green for Good usability (SUS ¿ 71), and dark green for Excellent usability (SUS ¿ 86).

### v4

@@ 2.1.3. Interface Design and the Obscuring of Epistemic Difference
2.2. Prior Work Targeting Cross-Lingual Knowledge Dissemination in Wikipedia
2.2.1. Surfacing vs. Consolidating Cross-Linguistic Differences
Prior work on cross-linguistic variation in Wikipedia has generally followed two contrasting orientations. One line of work treats differences across language editions primarily as obstacles to information access, motivating efforts to consolidate multilingual knowledge into a single, unified representation (Duh et al., 2013; Adar et al., 2009). While such approaches aim to improve coverage and navigability, they implicitly assume that cross-linguistic variation should be minimized, and risk obscuring the distinct perspectives embedded in distributed knowledge production by enforcing uniform representations (Schmidt and Bannon, 1992).
In contrast, a second orientation emphasizes surfacing cross-linguistic differences as meaningful expressions of positionality (Haraway, 2013). From this perspective, variation across language editions is not a shortcoming to be resolved, but an expected and valuable outcome of situated knowledge production. This framing aligns closely with our earlier discussion of epistemic equity, which calls for collaborative systems to recognize and sustain diverse ways of knowing rather than subsume them under a dominant perspective.
When translated into design terms, this orientation corresponds to an asset-based approach (Wong-Villacres et al., 2020; Kretzman and McKnight, 1993; Hardy and Thebault-Spieker, 2024). Rather than treating heterogeneity as a problem to be corrected, asset-based design views existing differences as resources for sense-making and learning. Applied to multilingual Wikipedia, this suggests that interfaces should help readers perceive and engage with cross-linguistic differences, instead of rendering them invisible through consolidation or assumed equivalence. Building on this perspective, our work extends prior efforts by exploring how interface design can surface positionality within everyday Wikipedia browsing practices.
@@ 2.2. Prior Work Targeting Cross-Lingual Knowledge Dissemination in Wikipedia
2.2.2. Prior Multilingual Interface Designs on Surfacing Cross-lingual Knowledge Differences.
Several systems have attempted to move beyond Wikipedia’s default ILLs by providing interfaces that surface content from multiple language editions. Manypedia (Massa and Scrinzi, 2012) displays two language editions side-by-side using machine translation, but it does not explicitly highlight factual inconsistencies or information gaps between them. WikiCompare (Roy et al., 2022) takes a step further by identifying topic-level differences. When a user reads an article in one language (e.g., Hindi), the interface highlights subtopics that are discussed in another language edition (e.g., English) but missing in the current version; clicking a highlight redirects the user to the corresponding English section. While informative, WikiCompare still operates at the topic level, rather than identifying specific factual discrepancies across languages. Moreover, neither Manypedia nor WikiCompare formally evaluates how such cross-lingual surfacing affects readers’ comprehension, behavior, or information-gathering efficacy.
The most ambitious multilingual interface to date is Omnipedia (Bao et al., 2012), which visualizes how different language communities conceptualize a topic by aggregating anchor-link structures across editions. This design allows simultaneous comparison of multiple languages but comes at the cost of requiring users to adopt an entirely new interaction paradigm. Users must access Omnipedia through its dedicated search interface, which contrasts with how most people encounter Wikipedia—via search engines and in-page browsing (Vincent and Hecht, 2021). As a result, Omnipedia’s workflow is difficult to integrate unobtrusively into readers’ existing habits on Wikipedia.
Complementing such reader-centered systems, the Wikipedia Diversity Observatory (WDO) (Miquel-Ribé and Laniado, 2020) offers editor-facing dashboards that quantify cultural and geographic content gaps across language editions and suggest missing but culturally relevant articles. However, WDO also operates outside Wikipedia’s native reading and editing interfaces and focuses only on article-level differences—an even coarser granularity than tools such as WikiCompare.
In summary, prior systems demonstrate the value of surfacing cross-lingual heterogeneity, yet they share several infrastructural limitations that restrict their practical impact: (1) they primarily expose topic-level rather than sentence-level differences; (2) they rely on workflows external to the natural reading experience, limiting their ability to integrate multilingual knowledge unobtrusively, As Grudin’s classic analysis of groupware failures highlights (Grudin, 1988), systems that impose new routines or additional work misaligned with existing practices frequently struggle to achieve sustained adoption; and (3) with the exception of Omnipedia, they lack empirical evaluation of how such designs actually affect readers.
2.2.3. Advances in multilingual factual knowledge retrieval
One of the challenges faced by prior systems that work at the article or topic-level was lack of access to computational tools that can recognize finer-grained differences between articles. Recent advances in multilingual NLP enable automatic detection of sentence-level factual gaps, which we leverage for developing our user-centric WikiGap design. Specifically, we rely on InfoGap (Samir et al., 2024, Sec. 3.2). The InfoGap paper corroborated through a large-scale analysis that English Wikipedia lacks facts from other language editions, including facts that may have a broader appeal. For example, Apple CEO Tim Cook’s association with the Russia-Ukraine war is only mentioned on Russian Wikipedia, but is conceivably noteworthy to the wider array of readers that engage with English Wikipedia. Building on InfoGap, our WikiGap system embeds those unique facts within Wikipedia’s native interface, overcoming limitations from the aforementioned interfaces.
@@ 3. WikiGap System
The goal of WikiGap is to challenge the English-as-superset assumption by making complementary knowledge from other language editions visible to readers during everyday Wikipedia use. Rather than treating multilingual differences as gaps to be corrected, WikiGap is designed to surface them in context, allowing readers to encounter cross-lingual knowledge while browsing English Wikipedia pages. We combine a computational pipeline that identifies sentence-level knowledge gaps across languages with a reader-facing interface that presents these differences through in-situ annotations and structured fact cards.
We begin with outlining the design requirements informed by user needs and foundational theories that shaped the WikiGap interface (§3.1). We then step behind the scenes to introduce InfoGap (Samir et al., 2024), which we use to detect language-exclusive facts lacking in English Wikipedia (§3.2). Furthermore, we describe how we extended and adapted InfoGap to detect and deliver language-exclusive facts, including how we filter, translate, and bind them to UI components at runtime (§3.3). We conclude this section by outlining the system implementation details (§3.4).
@@ 3.1.1. Design Requirements Informed by User Needs and Theory
Guided by an asset-based perspective that shifts the focus from correcting gaps to making differences visible and interpretable, we combined findings from our exploratory interviews with established theories of reading and information seeking to define the design requirements for WikiGap. These requirements share an overarching goal: enabling readers to access cross-lingual knowledge while preserving the familiar experience of browsing English Wikipedia. The findings from the exploratory interview, requirements, and design elements are summarized in Figure 2 and detailed below.
@@
We conducted exploratory interviews with four Computer Science graduate students (2 women, 2 men), all regular Wikipedia users; three spoke an additional language (Korean, Russian, or Hindi). Each 30-minute session followed an unstructured, conversational format guided by open-ended questions about their multilingual Wikipedia usage, perceptions of the Wikipedia ILL system, and ideas for improving cross-lingual content presentation. Participants reviewed an English Wikipedia article on mooncake–a topic unfamiliar to them–and reflected on how multilingual facts could be surfaced in this context. They proposed interface features to better support multilingual reading habits. We took notes during the sessions and synthesized them using thematic coding to derive the following design requirements for WikiGap.
@@
This requirement directly addresses participants’ tendency to overlook multilingual content due to the high cost of switching language editions. Participants expressed a strong desire to access multilingual information without breaking their reading experience. They reported rarely using Wikipedia’s ILLs (F1), noting the disorienting nature of full-page switches: “Even if I click the other language, I don’t know where to look for the thing I care about. It’s like starting over.” Such page jumps impose high navigation costs (Pirolli and Card, 1999), requiring users to reorient to a new layout and language, which discourages exploration.
@@
R2: Ensure clear language attribution and traceability of multilingual facts to support trust and credibility.
This requirement reflects readers’ need to assess the provenance of multilingual facts, especially when such information is absent from the English article. In our interviews, a recurring concern among participants was the trustworthiness of information surfaced from other language editions–particularly when such facts were not present in the English version. Participants emphasized that they wanted to know “where the information comes from,“ including which language edition it originated from and the ability to read it in its original context (F3), because it is crucial for accessing information’s credibility. This requirement highlights the importance of traceability and attribution in multilingual information access. Readers must be able to understand both the linguistic source and the original context of each surfaced fact. This entails more than just showing a translation, but rather also requires visual and interactive mechanisms that reinforce provenance and credibility.
@@
This requirement reflects the need to support different levels of engagement, recognizing that readers vary in how much cross-lingual information they want to encounter. User feedback revealed a range of interaction preferences: users want flexible access modes–both passive discovery and targeted retrieval (F4). Different participants had different expectations for how cross-lingual facts should be surfaced. Some favored passive discovery, where the system automatically highlights interesting differences without requiring user input. Others preferred more active control, such as being able to search for specific content or filter information by language or topic. They discussed concern over potential information overload if too many facts were presented at once.
@@ 3.1.2. Core Design Elements
This design addresses R1 (non-disruptive presentation) and R2 (traceability) to make multilingual differences visible at the moment of reading without interrupting readers’ flow, we implemented a subtle underlining system that signals the presence of additional information from other language editions. Each marker is country-color coded and appears as a patterned underline within the English article. When clicked, the side panel retrieves the fact relevant to that sentence, where the extension identifies complementary or missing information from another language edition. The underline then becomes more prominent, providing interaction feedback and inviting further exploration. Each language is associated with a distinct color – red for Chinese, blue for French, and green for Russian. This color coding is applied across both in-text highlights and corresponding elements in the side panel, reinforcing source attribution as emphasized in R2. As we elaborate in Sec. 3.3.2, only significant knowledge gaps identified by InfoGap are highlighted, rather than minor variations or differences in wording. This selective highlighting prevents information overload and ensures that users’ attention is drawn to substantive cross-lingual differences.
@@
To fulfill R1 and R2, we also developed a margin-anchored sidebar that displays the facts, translated to English and organized by their source language. When a user interacts with an in-text highlight, the corresponding fact card is revealed in the sidebar, which appears on the right side of the article. The sidebar is collapsible and can be pinned for extended browsing or hidden entirely to maintain a minimalist reading view. Each language section presents its facts in dedicated groups. Each fact is shown in English, accompanied by attribution information, including the source language tag and link to the original article. The close spatial relationship between in-text cues and sidebar content supports R1’s emphasis on minimizing reading disruption.
@@
At the heart of WikiGap’s cross-lingual surfacing system is the fact card interface, which presents individual facts from other language editions in a compact, standardized format. Each fact card includes: (1) the target fact, translated into English, (2) a color-coded badge indicating the source language, and (3) a hyperlink labeled “View on [Language] Wikipedia” that opens the source article in a new tab with the original sentence highlighted. This design primarily addresses R2 (transparent attribution) by balancing at-a-glance comprehension with direct access to original context.
@@ 3.2. The InfoGap Pipeline
While the interface makes multilingual differences visible to readers, InfoGap provides the backend capability that enables such differences to be identified at sentence-level granularity. It was a previously developed computational pipeline introduced by Samir et al. (Samir et al., 2024), which detects factual misalignments between Wikipedia articles across languages. This subsection introduces the core components of the original InfoGap system, which we use to identify sentence-level knowledge gaps between English and other language editions. In the following section, we will describe how we extended this pipeline to support additional languages and adapted its outputs for integration into the WikiGap interface.
@@ 3.3. Extending and Integrating InfoGap for WikiGap
3.3.1. Rationale for Language Selection
In this study, we intentionally focused on English and three additional language editions, namely Russian, French, and Chinese. We chose the first two because the original InfoGap pipeline was developed for detecting cross-lingual gaps in biographical articles between English and either Russian or French. We included them because of they have been validated in the research to ensure accuracy. To generalize the method to a new domain and an additional language, we incorporated support for Chinese into the fact decomposition and verification modules, as some co-authors were proficient in Chinese. This involved custom preprocessing for tokenization, crafting prompts, and conducting human evaluation on Chinese texts to confirm that InfoGap performs comparably to its performance on the original languages studied by Samir et al. (2024).
@@ 4. Evaluation Methods
To examine whether surfacing cross-lingual knowledge gaps within the English Wikipedia interface can meaningfully support multilingual engagement, we conducted a mixed-methods user study combining quantitative performance measures with qualitative feedback. The study assesses usability, knowledge acquisition, and users’ perceptions of the system and the cross-lingual content it surfaces. Table 1 summarizes the study metrics and analysis. [^3] The study protocol received approval from our institutional REB.
@@ 4.1. Participants
The study followed a within-subject design, in which each participant was assigned two different topics out of the total five topics (Sec. 4.3). Each participant read one article using the WikiGap extension and another article using the default Wikipedia interface (with ILLs) as a control. We selected interlanguage links (ILLs) as the control to reflect the default, in-context Wikipedia experience and to simulate an extension built on top of existing knowledge-seeking workflows; moreover, prior multilingual interface designs are not publicly available for deployment. The assignment of treatment and control conditions was randomized for each participant.
@@ 4.3. Materials and Topic Assignment
Each participant is assigned two distinct topics, one for each condition (control and WikiGap). Topic assignment was designed to minimize cultural confounds and familiarity‑driven performance differences. Specifically, the food topic chosen for each participant could not be strongly linked to the participant’s self‑reported cultural or ethnic background. To reduce individual-performance bias, each topic was assigned in every condition to exactly four different participants.
@@ 5. Findings
We report findings from our mixed-method evaluation of WikiGap. Together, these findings show how making multilingual knowledge differences visible within the English Wikipedia interface affects users’ performance, preferences, and engagement with non-English content. Across quantitative measures and qualitative feedback, participants consistently described shifts in how they accessed information, perceived Wikipedia’s completeness, and engaged with multilingual knowledge during routine fact-finding.
1. Making Multilingual Differences Visible Improves User Preference and Reading Experience (D1, D2, D3)
We evaluated user preference using the System Usability Scale (SUS). The average SUS scores for each condition are presented in Figure 6, along with general SUS cutoff scores for fair, good, and excellent usability (Bangor et al., 2009). Overall, WikiGap achieved a substantially higher usability score compared to the control. WikiGap falls within the “excellent” usability category, while the default Wikipedia page with ILL scored below the “fair” threshold. A paired-samples t-test revealed a significant difference in scores between the WikiGap condition (M = 82.5, SD = 10.9) and the control condition (M = 48.2, SD = 18.6), t(20) = 6.60, p ¡ .001. (Figure 6).
Figure 6. Box plot showing the System Usability Scale (SUS) scores for each condition. Blue represents the control condition (no WikiGap), and orange represents the treatment condition (with WikiGap). Horizontal dashed lines represent standard usability benchmarks in varying shades of green: light green for Fair usability (SUS ¿ 51), medium green for Good usability (SUS ¿ 71), and dark green for Excellent usability (SUS ¿ 86).